Sean's bank charges him $9.75 per month for the first 10 transactions and $1.35 for each additional transaction. Calculate Sean's fee if he made the following number of transactions: a) 9 b) 21
Answer by stanbon(75887)

Sean's bank charges him $9.75 per month for the first 10 transactions and $1.35 for each additional transaction. Calculate Sean's fee if he made the following number of transactions:

a) 9
F(9) = 9.75*9 = $87.75
-----------------------
b) 21
F(21) = 10*9.75 + 11*1.35 = $112.57
---------------
Cheers,
Stan H.
-------------